OpenStudy (daryan):

integral 1/((3)+(sqrt(x+2)) please show steps

OpenStudy (miamoorz):

can you please write it in a better form? with the equation editor or upload a picture?

OpenStudy (miamoorz):

\[\int\limits_{}^{} (\frac{ 1 }{ 3+\sqrt{x+2} }) dx\] is it like this?

OpenStudy (daryan):

yes

OpenStudy (miamoorz):

do you just want to integrate it? or what? i mean what does the question says?

OpenStudy (daryan):

just integrate

OpenStudy (miamoorz):

Do you have a final answer for that i just wanna make sure?

OpenStudy (daryan):

answer is 2( sqrt(x+2) - 3log(sqrt(x+2)))

OpenStudy (marcelie):

u substitution hmm

OpenStudy (photon336):

yeah I think you can use u substitution here.

OpenStudy (astrophysics):

What if you multiply by the conjugate :O

OpenStudy (astrophysics):

Nvm you can use 2 u subs

OpenStudy (phi):

***answer is 2( sqrt(x+2) - 3log(sqrt(x+2)))**** I get \[ 2 \left( \sqrt{x+2} - 3\ln(\sqrt{x+2}+3) \right)\] or \[ 2\sqrt{x+2} - 6\ln(\sqrt{x+2}+3) \] using the substitution \[ u = \sqrt{x+2} \\ du = \frac{1}{2} \frac{1}{\sqrt{x+2}} dx = \frac{dx}{2u} \\ dx = 2 u \ du \]